# Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark

### v2

@@ 2.3.1 Confidence-based approaches
Confidence-based approaches are widely used in the literature. The simplest trigger model of this kind consists in monitoring a quantity related to the confidence of the prediction over time and triggering class prediction as soon as a threshold value is exceeded. The confidence metric monitored can take different forms. For example, a baseline approach, referred to as Proba Threshold [^3] Baseline implemented in the aeon (?) library : https://urlz.fr/qmWl in the remainder of this paper, involves monitoring $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$ the highest conditional probability estimated by the classifier. This baseline example is qualified as instant-based method, since it takes as input only the last confidence score available at time $t$. Another type of approaches, qualified as sequence-based, monitors the entire sequence of past confidence scores, and triggering a prediction is made conditionally on a particular property of this sequence. Accordingly, trigger functions can either take as input a scalar value, e.g. $g(\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t}))$, in the case of instant based approaches, or a sequence of scalar values, e.g. $g(\{\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{\tau})\}_{1\leq\tau\leq t})$, in the case of sequence-based approaches (see Section 3.1).
2.3.2 cost-informed at testing time
Given that an ECTS approach will ultimately be evaluated on the average cost of using it (see Equation 4), it seems natural to exploit the cost values at testing time, in order to trigger predictions at optimal moments. Methods such as Economy (?) and 2step/NoCluster (?) do that. They can thus be qualified as “cost-informed at testing time”.
Other approaches use instead trigger functions that, once learned, do not take into account the cost at testing time, but rely on other measures such as, for instance, the confidence of the prediction. This is the case of the SR approach (?). Therefore, these approaches can be qualified as “cost-uninformed at testing time”.
Notice that some approaches are cost-informed during training but not during inference. This is the case with the SR approach, which is cost-informed at training time since it uses costs to optimize its parameters, and cost-uninformed at testing time since the resulting trigger function does not use cost values during inference. Table 1 shows these two different properties for each approach in the literature.
2.3.3 Blind, Myopic and Anticipation-based decisions
Some separable approaches consider the output of the classifier $h$ at time $t$ to decide whether this is the right time to make a prediction. For instance, stopping the process when the confidence in the classification $h_{t}(\mathbf{x}_{t})$ is above some threshold, or when the difference of confidence between the best two predictions exceeds some value. These methods can be described as myopic since they only look at the current time step $t$, without trying to guess the future.
But there is a another possibility. As was first noted by ? (?), the ECTS problem can be cast as a LUPI (Learning Using Privileged Information) problem (?).
@@ 2.3.3 Blind, Myopic and Anticipation-based decisions
In this scenario, the learner can benefit at the training time of privileged information that will not be available at test time. Formally, the training set can be expressed as $\mathcal{T}=\{(\mathbf{x}_{i},\mathbf{x}_{i}^{\star},y_{i})\}$, where $\mathbf{x}_{i}$ is what is observable and $\mathbf{x}_{i}^{\star}$ is some additional information not available when the prediction must be made. This is exactly what happens in the ECTS problem. Whereas at test time, only $\mathbf{x}_{t}$ is available, during training the complete time series are known. This brings the possibility to learn what are the likely future of an incoming time series $\mathbf{x}_{t}$ provided it comes from the same distribution. Hence, it becomes also possible to guess the cost to be optimized for all future time steps, and therefore to wait until the moment seems the best. This type of approach can be said anticipation-based (also called non-myopic in the literature). Because more information from the training set is exploited, it can be expected that these methods outperform myopic and blind ones.
@@ 2.4 Choice of the classification component
One source of difficulty when devising an ECTS method in the separable setting is that inputs differ from one time step to another. The number of measurements, and hence the input dimension, varies. Two approaches have been used to deal with the problem.
@@
    A set of classifiers $\{h_{t}\}_{t\in[1,T]}$ is learned, each dedicated to a given time step $t$, and thus a given input dimension. In practice, authors often chose a limited subset of timestamps, usually a set of twenty (one measurement every 5% of the length of the time series), to restrict the number of classifiers to learn and therefore the associated computational cost.
@@
    A single classifier $h$ is used for all possible incoming time series $\mathbf{x}_{t}$. One way of doing this is to “project” an input $\mathbf{x}_{t}$ of dimension $t\times d$, if $d$ is the dimension of an observation at time $t$ (i.e. multi-valued time series), into a fixed dimensional vector whatever $t$ and $d$. This may simply be the mean value and standard deviation of the available measurements (multiplied by the dimension $d$) or the result of a more sophisticated feature engineering as tested by ? (?). Deep learning architectures can also be used to learn an encoding of the time series in an intermediate layer. For instance, ? (?) use a CNN architecture, and ? (?) a FCN one. ? (?) show that using deep neural architectures often perform well for time series classification.
@@
3 State of the art
In this section, we position various existing proposed methods along the dimensions of the taxonomy presented in Section 2 (see Table 1). Then, we examine the methods and how they exemplify solutions to the general questions raised in the taxonomy. Thus, while each method combines solutions to all questions, in the following, we underline how each one brings an original solution to one specific problem. For instance, the Economy approach is separable, anticipation-based, and cost-informed at testing time. However, we’re focusing here on the anticipation-based dimension, since this is one of the first methods to have emphasized it and introduced an original solution for this aspect.
References Classifier(s) (collection ✓) End2end Confidence Anticipation Cost informed
EDSC (?) Shapelet
ECTS’ (?) 1NN
Reject (?) SVM
RelClass (?) QDA, Linear SVM
[…]
2step/NoCluster (?) Linear SVM (✓) train & test
ECDIRE (?) Gausian Process (✓)
Stopping Rule (?) Gaussian Process (✓) train
[…]
ECEC (?) WEASEL (✓) train
DDQN (?) MLP train
TEASER (?) WEASEL (✓) train
[…]
DETSCNet (?) TCN train
CALIMERA (?) MiniROCKET (✓) train
ELECTS (?) LSTM train
[…]
EarlyStop-RL (?) MLP train
Table 1: Table of published methods for the ECTS problem with their properties along dimensions underlined in the taxonomy.
@@ 3.1 Confidence-based approaches
Most ECTS methods to date are separable, confidence-based, cost-informed at training time, and are not anticipation-based. They implement separately the prediction and the triggering components, they learn them using the costs, hence they are cost-informed at training time, but they decide to trigger a decision based on the information available at the current time step $t$ without trying to anticipate the likely future, and they base their decision upon the confidence of the predictions made by the classifier.
@@ 3.1.1 Instant-based decision criterion
$\bullet$ One basic method is to monitor $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$, the highest conditional probability estimated by the classifier, which is a simple measure of classifier confidence over time. As soon as it exceeds a value, which is an hyper parameter of the method, a prediction is made. We call this method Proba Threshold and use it as baseline for comparison later in our experiments.
$\bullet$ The Reject method (?) uses ensemble consensus as a confidence measure. For each time step, first (i), a pool of classifiers is trained by varying their hyperparameters (i.e. SVMs); then (ii), the most accurate of these are selected; and (iii) the pair of classifiers minimizing their agreement in predictions is chosen to form the ensemble. Finally, the prediction is triggered as soon as both classifiers in the ensemble predict the same class value. In this case, the monitored confidence measure is binary (agreement or disagreement), there is no trigger threshold and thus this trigger model is free of hyper-parameters [^4] The Reject approach involves choosing the number of classifiers trained in step (i) and selected in step (ii) that could be considered as hyper-parameters of the monitored confidence measure..
$\bullet$ Hidden Markov Models (HMMs) are naturally suited to the classification of online sequences. An HMM is learned for each class, and at each time step $t$, the class to be preferred is the one with the highest a posteriori probability given $\mathbf{x}_{t}$. However, the decision to make a prediction now or to postpone it must then involve a threshold so that the prediction is only made if the a posteriori probability of the best HMM is sufficiently high or is greater than that of the second-best. In reaction to this, ? (?) propose to replace the standard HMM with imprecise HMMs based on the concept of credal classifier. This eliminates the need to choose a threshold, since a decision is made when one classification “dominates” (according to a criterion based on probability intervals) all the others.
@@
This is one dimension used in the Stopping Rule (SR) approach (?). Specifically, the output of the system is defined as:
$g(h({\mathbf{x}}_{t}))=\begin{cases}\emptyset&\mbox{if extra measures are queried;}\\ \hat{y}=\operatorname*{arg\,max}_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})&\mbox{when }\gamma_{1}\,p_{1}+\gamma_{2}\,p_{2}+\gamma_{3}\,\frac{t}{T}>0\end{cases}$ (9)
where $p_{1}$ is the largest posterior probability $p(y|{\mathbf{x}}_{t})$ estimated by the classifier $h$, $p_{2}$ is the difference between the two largest posterior probabilities, and $\frac{t}{T}$ represents the proportion of the incoming time series at time $t$. The parameters $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$ are learned from the training set.
$\bullet$ Using the same notations as SR, the Early Classification framework based on class DIscriminativeness and RELiability (Ecdire) (?) finds the earliest timestamp for which a threshold applied on $p_{1}$ is reached (defined as in Equation 9). Then, the quantity $p_{2}$ is monitored, and a second threshold is applied to trigger the prediction.
$\bullet$ A different class of methods relies on searching telltale representations of subsequences, such that if the incoming time sequence $\mathbf{x}_{t}$ matches one or more of these representations, then its class can be predicted. Typically, these representations take the form of shapelets that discriminate well one class from the others (?). For instance, the Early Distinctive Shapelet Classification (Edsc) method learns a distance threshold for each shapelet, based on the computation of the Euclidean distance between the considered subsequence and all others valid subsequences in the training set (?). It selects a subset of them, based on a utility measure that combines precision and recall, weighted by the earliness. A prediction is made as soon as $\mathbf{x}_{t}$ matches one of these shapelets well enough. Because this family of methods is computationally expensive, extensions have been developed to reduce the computational load (?, ?). Other extensions aimed at improving the reliability of the predictions (?, ?), and tackling multivariate time series (?, ?, ?, ?).
@@ 3.1.2 Sequence-based decision criterion
$\bullet$ The Effective Confidence-based Early Classification (Ecec) (?) proposes a confidence measure based on the sequence of predicted class values, from the first one observed to the current timestamp. At each time step, this approach exploits the precision of the classifier to estimate the probability for each possible class value $y\in{\cal Y}$ of being correct if predicted. Then, assuming that successive class predictions are independent, the proposed confidence measure represents the probability that the last class prediction is correct given the sequence of predicted class values. The proposed confidence measure is monitored over time, and prediction is triggered if this measure exceeds a certain threshold $\gamma$ tuned as the single hyper-parameter.
$\bullet$ The Teaser (Two-tier Early and Accurate Series classifiER) (?) approach considers the problem of whether or not a prediction should be triggered as a classification task, the aim of which is to discriminate between correct and bad class predictions. As the authors point out, the balance of this classification task varies according to the time step considered $t\in[1,T]$. Indeed, assuming there is an information gain over time, there are fewer and fewer bad decisions as new measurements are received (or even no bad decisions after a while, i.e. $\forall\;t>t^{\prime}\;(0<t^{\prime}\leq T)$ for some datasets). To exploit this idea, a collection of one-class SVMs is used, learning hyper-spheres around the correct predictions for each time step. A prediction is triggered when it falls within these hyper-spheres for $\nu$ consecutive time steps ($\nu$ being a parameter of the method).
$\bullet$ The Second-Order Confidence Network approach (Socn) (?) considers, as does Teaser, the same classification task aiming to discriminate between correct and bad predictions. To learn this task, a transformer (?) is used, taking as input the complete sequence of conditional probabilities estimated by the classifier $h$, from the first time step, up to the current time step. A confidence threshold $\nu$ is learned by minimizing the same cost function as (?), above which the prediction is considered reliable and therefore triggered.
@@ 3.2 Anticipation-based methods
One way of designing approaches that anticipate future measurements is to achieve classification of an incomplete time series while guaranteeing a minimum probability threshold according to which the same decision would be made on the complete series. This is the case of the Reliability Classification (RelClass) approach (?). Assuming that the measurements are i.i.d. and generated by a Gaussian process, this approach estimates $p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})$ the conditional probability of the entire time series ${\mathbf{x}}_{T}$ given an incomplete realization ${\mathbf{x}}_{t}$ and thus derives guarantees of the form:
[…]
where ${\mathbf{x}}_{T}$ is a random variable associated with the complete times series, $\gamma$ is a confidence threshold, and $h_{T}$ is the classifier learned over complete times series. At each time step $t$, $p(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t})$ is evaluated and a prediction is triggered if this term becomes greater than the threshold $\gamma$, which is the only hyper-parameter to be tuned.
Another way of implementing anticipation-based approaches is to exploit the continuations of training time series, which are full-length. One of the first method for ECTS (?) has been derived into such an anticipation-based approach (?). The first, called Early Classification on Time Series (Ects) (?), exploits the concept of Minimum Prediction Length (MPL), defined as the earliest time step for which the predicted label should not change for the incoming time series $\mathbf{x}_{t}$ from $t$ to $T$. This is estimated by looking for the 1NN of $\mathbf{x}_{t}$ in the training set, and check whether from $t$ onward, its predicted label did not change. To be more robust, the MPL is defined based on clusters computed on full-length training time series to estimate the best decision time. The approach has been extended later on to speed up the learning stage (?). This method looks in its own way at the likely future of $\mathbf{x}_{t}$ i.e. an incomplete time series belongs to a cluster whose continuations are known - and thus can be considered as an anticipation-based method.
? (?) present a method that claims explicitly to be “non myopic” in that a decision is taken at time $t$ only insofar as it seems that no better time for prediction is to be expected in the future. In order to do this, the family of Economy methods estimates the future cost expectation based on the incoming time series $\mathbf{x}_{t}$. This can be done since the training data consists of full-length time series and therefore a Learning Using Privileged Information (LUPI) (?) is possible.
More formaly, the objective is to trigger a decision when $\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]$ is minimal, with:
@@
A key challenge in this framework is to design approaches achieving the most useful partition for predicting decision costs expectation. In the first article which presents this framework (?), a method, called Economy-$K$, is designed as follows. (i) A partition of training examples is first performed by a K-means algorithm ; (ii) then a simple model uses the Euclidean distance as a proxy of the probability that $\mathbf{x}_{t}$ belongs to each group; (iii) the continuation of training time series within each group is exploited to predict the cost expectation for future time steps.
@@
In order to avoid the clustering step with the associated choice of hyper parameters, (?) presented a variant called NoCluster which uses the 1 nearest neighbor in the training set in order to guess the likely future of $\mathbf{x}_{t}$.
@@
Then, Economy-$\gamma$ was introduced in (?) which relies on a supervised method to define a confidence-based partition of training time series. The algorithm, dedicated to binary classification problems, is designed as follows: (i) a collection of partitions is constructed by discretizing the output of each classifier $\{h_{i}\}_{i\in[1,T]}$ into equal-frequency intervals, the groups thus formed correspond to confidence levels for each time step; (ii) at the current time $t$, the incoming time series $\mathbf{x}_{t}$ belongs to only one group, since the output of the classifier $h_{t}$ falls within a particular confidence level; (iii) then, a Markov chain model is trained to estimate the probabilities of the future time step confidence levels. Economy-$\gamma$-max (?) generalizes this approach to multi-class problems, aggregating the multiple conditional probabilities in the classifiers’ output by using only the most probable class value.

### v6

@@ 2.2.2 Confidence-based approaches
The simplest trigger model of this kind consists in monitoring a quantity related to the confidence of the prediction over time and triggering class prediction as soon as a threshold value is exceeded. The confidence metric monitored can take different forms. For example, a baseline approach, referred to as Proba Threshold [^3] Baseline implemented in the aeon (Middlehurst et al., 2024a) library : https://urlz.fr/qmWl in the remainder of this paper, involves monitoring $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$ the highest conditional probability estimated by the classifier. This baseline example is qualified as instant-based method, since it takes as input only the last confidence score available at time $t$. Another type of approaches, qualified as sequence-based, monitors the entire sequence of past confidence scores, and a triggering prediction is made conditionally on a particular property of this sequence. Accordingly, trigger functions can either take as input a scalar value, e.g. $g(\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t}))$, in the case of instant-based approaches, or a sequence of scalar values, e.g. $g(\{\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{\tau})\}_{1\leq\tau\leq t})$, in the case of sequence-based approaches (see Section 3.1).
Those kind of examples can be described as myopic since they only look at the current time step $t$ (or past ones as well in the sequence-based case), without trying to anticipate the likely future, i.e. not directly having a decision made based on a forecast of some form. The anticipation-based approaches do this.
2.2.3 Anticipation-based decisions
@@
As was first noted by Achenchabe et al. (2021), the ECTS problem can be cast as a LUPI (Learning Using Privileged Information) problem (Vapnik & Vashist, 2009). In this scenario, the learner can benefit at the training time from privileged information that will not be available at test time. Formally, the training set can be expressed as $\mathcal{T}=\{(\mathbf{x}_{i},\mathbf{x}_{i}^{\star},y_{i})\}$, where $\mathbf{x}_{i}$ is what is observable and $\mathbf{x}_{i}^{\star}$ is some additional information not available when the prediction must be made. This is exactly what happens in the ECTS problem. Whereas at test time, only $\mathbf{x}_{t}$ is available, during training the complete time series is known. This brings the possibility to learn what are the likely futures of an incoming time series $\mathbf{x}_{t}$ provided it comes from the same distribution. Hence, it becomes also possible to guess the cost to be optimized for all future time steps, and therefore to wait until the moment seems the best. This type of approach can be said anticipation-based (also called non-myopic in the literature). They can come with many flavors, as long as their decisions are based on some kind of anticipation (see Section 3.2). Because more information from the training set is exploited, it can be expected that these methods outperform myopic ones.
@@ 2.3 Choice of the classification component
[…]
One source of difficulty when devising an ECTS method in the separable setting is that, at testing time, inputs differ from one time step to another. When an incoming time series is progressively observed, the number of measurements, and hence the input dimension, varies. Two approaches have been used to deal with the problem.
@@
    A set of classifiers $\{h_{t}\}_{t\in[1,T]}$ is learned, each dedicated to a given time step $t$, and thus a given input dimension. In practice, authors often choose a limited subset of timestamps, usually a set of twenty (one measurement every 5% of the length of the time series), to restrict the number of classifiers to learn and therefore the associated computational cost.
@@
    A single classifier $h$ is used for all possible incoming time series $\mathbf{x}_{t}$. One way of doing this is to “project” an input $\mathbf{x}_{t}$ of dimension $t\times d$, if $d$ is the dimension of an observation at time $t$ (i.e. multi-valued time series), into a fixed dimensional vector whatever $t$ and $d$. This may simply be the mean value and standard deviation of the available measurements (multiplied by the dimension $d$) or the result of a more sophisticated feature engineering as tested by Skakun et al. (2017). Deep learning architectures can also be used to learn an encoding of the time series in an intermediate layer (Wang et al., 2016; Sawada et al., 2022).
@@
3 An organized state-of-the-art of separable methods
In this section, approaches from the literature are considered and organized around two key notions from the introduced taxonomy: confidence-based and anticipation-based.
Subsection 3.1 presents methods whose decisions are triggered in a myopic way, based on some confidence measure. Subsection 3.2 describes approaches using a non-myopic decision criterion, which attempts to anticipate likely continuations. Table 1 shows how various approaches described in the literature can be organized using the characteristics underlined in the proposed taxonomy.
Table 1: Table of published separable methods for the ECTS problem with their properties along dimensions underlined in the taxonomy. Note that void values indicate that the corresponding property is not present in the referred system.
References Classifier(s) (collection ✓) Confidence Anticipation Cost informed
Reject (Hatami & Chira, 2013) SVM instant
iHMM (Antonucci et al., 2015) HMM instant
ECDIRE (Mori et al., 2017b) Gaussian Process (✓) instant
Stopping Rule (Mori et al., 2017a) Gaussian Process (✓) instant
ECEC (Lv et al., 2019) WEASEL (✓) sequence
TEASER (Schäfer & Leser, 2020) WEASEL (✓) sequence
SOCN (Lv et al., 2023) FCN sequence
ECTS (Xing et al., 2012) 1NN instant
RelClass (Parrish et al., 2013) QDA, Linear SVM instant
2step/NoCluster (Tavenard & Malinowski, 2016) Linear SVM (✓) ✓(test)
ECONOMY-$\gamma$-max (Zafar et al., 2021) XGBoost ++ tsfel (✓) ✓(test)
CALIMERA (Bilski & Jastrzebska, 2023) MiniROCKET (✓)
FIRMBOUND (Ebihara et al., 2025) LSTM
@@ 3.1.1 Instant-based decision criterion
$\bullet$ One basic method is to monitor $\max_{y\in{\cal Y}}p(y|{\mathbf{x}}_{t})$, the highest conditional probability estimated by the classifier, which is a simple measure of classifier confidence over time. As soon as it exceeds a value, which is a hyperparameter of the method, a prediction is made. We call this method Proba Threshold and use it as a baseline for comparison later in our experiments.
$\bullet$ The Reject method (Hatami & Chira, 2013) uses ensemble consensus as a confidence measure. For each time step, first (i), a pool of classifiers is trained by varying their hyperparameters (i.e. SVMs); then (ii), the most accurate of these are selected; and (iii) the pair of classifiers minimizing their agreement in predictions is chosen to form the ensemble. Finally, the prediction is triggered as soon as both classifiers in the ensemble predict the same class value.
$\bullet$ Hidden Markov Models (HMMs) are naturally suited to the classification of online sequences. An HMM is learned for each class, and at each time step $t$, the class to be preferred is the one with the highest a posteriori probability given $\mathbf{x}_{t}$. However, the decision to make a prediction now or to postpone it must then involve a threshold so that the prediction is only made if the a posteriori probability of the best HMM is sufficiently high or is greater than that of the second-best. In reaction to this, Antonucci et al. (2015) propose to replace the standard HMM with imprecise HMMs based on the concept of credal classification. This eliminates the need to choose a threshold, since a decision is made when one classification “dominates” (according to a criterion based on probability intervals) all the others.
@@
This is one dimension used in the Stopping Rule (SR) approach (Mori et al., 2017a). Specifically, the output of the system is defined as:
[…]
where $p_{1}$ is the largest posterior probability $p(y|{\mathbf{x}}_{t})$ estimated by the classifier $h$, $p_{2}$ is the difference between the two largest posterior probabilities, and $\frac{t}{T}$ represents the proportion of the incoming time series at time $t$. The parameters $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are learned from the training set, using genetic algorithm.
$\bullet$ Using the same notations as SR, the Early Classification framework based on class DIscriminativeness and RELiability (Ecdire) (Mori et al., 2017b) finds the earliest timestamp for which a threshold applied on $p_{1}$ is reached (defined as in Equation 5). Then, the quantity $p_{2}$ is monitored, and a second threshold is applied to trigger the prediction.
$\bullet$ Ringel et al. (2024) use the Learning Then Test (LTT) (Angelopoulos et al., 2021) calibration framework to address ECTS. In practice, the proposed approach greedily computes thresholds at each time step, in order to monitor some conditional control risk measure, given a pre-defined error rate. This paper investigates text applications.
$\bullet$ Historically, a related scenario predates the ECTS problem but is different. In the sequential decision making and optimal statistical decisions frameworks (DeGroot, 2005; Berger, 1985), the successive measurements are supposed to be independently and identically distributed (i.i.d.) according to a distribution of unknown “parameter” $\theta$. The problem is to determine as soon as possible whether the measurements have been generated by a distribution of parameter $\theta_{0}$ (hypothesis $H_{0}$) or of parameter $\theta_{1}$ (hypothesis $H_{1}$) with $\theta_{0}\neq\theta_{1}$. In the Wald’s Sequential Probability Ratio Test (Wald & Wolfowitz, 1948; Ghosh & Sen, 1991), the log-likelihood ratio $R_{t}\;=\;\log\frac{P(\langle x_{1}^{i},\ldots,x_{t}^{i}\rangle\;|\;y=-1)}{P(\langle x_{1}^{i},\ldots,x_{t}^{i}\rangle\;|\;y=+1)}$ is computed and compared with two thresholds that are set according to the required error of the first kind $\alpha$ (false positive error) and error of the second kind $\beta$ (false negative error). This beautiful setting allows one to get optimal decision times at the cost of being able to compute the log-likelihood. However, it differs from the ECTS problem, where successive observations are dependent. The i.i.d. assumption being not valid for the ECTS problem, a generalization to the non-i.i.d. case was proposed by Tartakovsky et al. (2014), providing guarantees for the asymptotic case (with $T\rightarrow\infty$). Despite this latter limitation, Ebihara et al. (2025) has recently applied this type of approach to ECTS with finite time horizons. The authors propose practical ways of both estimating $R_{t}$ (Ebihara et al., 2023) and triggering times by solving a backward induction problem (Tartakovsky et al., 2014).
@@ 3.1.2 Sequence-based decision criterion
$\bullet$ The Effective Confidence-based Early Classification (Ecec) (Lv et al., 2019) proposes a confidence measure based on the sequence of predicted class values, from the first one observed to the current timestamp. At each time step, this approach exploits the precision of the classifier to estimate the probability for each possible class value $y\in{\cal Y}$ of being correct if predicted. Then, assuming that successive class predictions are independent, the proposed confidence measure represents the probability that the last class prediction is correct given the sequence of predicted class values. The proposed confidence measure is monitored over time, and prediction is triggered if this measure exceeds a certain threshold $\gamma$ tuned as the single hyperparameter.
$\bullet$ The Teaser (Two-tier Early and Accurate Series classifiER) (Schäfer & Leser, 2020) approach considers the problem of whether or not a prediction should be triggered as a classification task, the aim of which is to discriminate between correct and bad class predictions. As the authors point out, the balance of this classification task varies according to the time step considered $t\in[1,T]$. Indeed, assuming there is an information gain over time, there are fewer and fewer bad decisions as new measurements are received (or even no bad decisions after a while, i.e. $\forall\;t>t^{\prime}\;(0<t^{\prime}\leq T)$ for some datasets). To exploit this idea, a collection of one-class SVMs is used, learning hyper-spheres around the correct predictions for each time step. A prediction is triggered when it falls within these hyper-spheres for $\nu$ consecutive time steps.
$\bullet$ The Second-Order Confidence Network approach (Socn) (Lv et al., 2023) considers, as does Teaser, the same classification task aiming to discriminate between correct and bad predictions. To learn this task, a transformer (Vaswani et al., 2017) is used, taking as input the complete sequence of conditional probabilities estimated by the classifier $h$, from the first time step, up to the current time step. A confidence threshold $\nu$ is learned by minimizing the same cost function as Lv et al. (2019) do, above which the prediction is considered reliable and therefore triggered.
@@ 3.2 Anticipation-based methods
$\bullet$ One way of designing approaches that anticipate future measurements is to achieve classification of an incomplete time series while guaranteeing a minimum probability threshold, according to which the same decision would be made on the complete series. This is the case of the Reliability Classification (RelClass) approach (Parrish et al., 2013). Assuming that the measurements are i.i.d. and generated by a Gaussian process, this approach estimates $p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})$ the conditional probability of the entire time series ${\mathbf{x}}_{T}$ given an incomplete realization ${\mathbf{x}}_{t}$ and thus derives guarantees of the form:
$p\bigl(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t}\bigr)\,=\,\int_{{\mathbf{x}}_{T}\text{ s.t. }h_{T}({\mathbf{x}}_{T})=y}\,p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})\,d{\mathbf{x}}_{T}\,\geq\,\gamma$
where ${\mathbf{x}}_{T}$ is a random variable associated with the complete time series, $\gamma$ is a confidence threshold, and $h_{T}$ is the classifier learned over the complete time series. At each time step $t$, $p(h_{T}({\mathbf{x}}_{T})=y|{\mathbf{x}}_{t})$ is evaluated and a prediction is triggered if this term becomes greater than the threshold $\gamma$, which is the only hyperparameter to be tuned.
$\bullet$ Another way of implementing anticipation-based approaches is to exploit the continuations of training time series, which are full-length. One of the first methods for ECTS has been derived into such an anticipation-based approach. The first, called Early Classification on Time Series (Ects) (Xing et al., 2009), exploits the concept of Minimum Prediction Length (MPL), defined as the earliest time step for which the predicted label should not change for the incoming time series $\mathbf{x}_{t}$ from $t$ to $T$. This is estimated by looking for the 1NN of $\mathbf{x}_{t}$ in the training set, and checks whether from $t$ onward, its predicted label does not change. To be more robust, the MPL is defined based on clusters computed on full-length training time series to estimate the best decision time. The approach has been extended later on to speed up the learning stage (Xing et al., 2012). This method looks in its own way at the likely future of $\mathbf{x}_{t}$ - i.e. an incomplete time series belongs to a cluster whose continuations are known - and thus can be considered as an anticipation-based method.
$\bullet$ Dachraoui et al. (2015) present a method that claims explicitly to be “non-myopic” in that a decision is taken at time $t$ only insofar as it seems that no better time for prediction is to be expected in the future. In order to do this, the family of Economy methods estimates the future cost expectation based on the incoming time series $\mathbf{x}_{t}$. This can be done since the training data consists of full-length time series and therefore a Learning Using Privileged Information (LUPI) (Vapnik & Vashist, 2009) is possible.
More formally, the objective is to trigger a decision when $\mathbb{E}_{y,\hat{y}}[\mathcal{L}(\hat{y},y,t)|\mathbf{x}_{t}]$ is minimal, with:
@@
A key challenge in this framework is to design approaches achieving the most useful partition for predicting decision costs expectation. In the first article which presents this framework Dachraoui et al. (2015), a method, called Economy-$K$, is designed as follows. (i) A partition of training examples is first performed by a K-means algorithm ; (ii) then a simple model uses the Euclidean distance as a proxy of the probability that $\mathbf{x}_{t}$ belongs to each group; (iii) the continuation of training time series within each group is exploited to predict the cost expectation for future time steps.
@@
In order to avoid the clustering step with the associated choice of hyperparameters (Tavenard & Malinowski, 2016) presented a variant called NoCluster which uses the 1-nearest neighbor in the training set in order to guess the likely future of $\mathbf{x}_{t}$.
@@
Then, Economy-$\gamma$ was introduced by Achenchabe et al. (2021) which relies on a supervised method to define a confidence-based partition of training time series. The algorithm, dedicated to binary classification problems, is designed as follows: (i) a collection of partitions is constructed by discretizing the output of each classifier $\{h_{i}\}_{i\in[1,T]}$ into equal-frequency intervals, the groups thus formed correspond to confidence levels for each time step; (ii) at the current time $t$, the incoming time series $\mathbf{x}_{t}$ belongs to only one group, since the output of the classifier $h_{t}$ falls within a particular confidence level; (iii) then, a Markov chain model is trained to estimate the probabilities of the future time step confidence levels. Economy-$\gamma$-max (Zafar et al., 2021) generalizes this approach to multi-class problems, aggregating the multiple conditional probabilities in the classifiers’ output by using only the most probable class value.